The drawing below shows a square with side a. A straight line intersects the square and encloses

an area A. The heights x and y on the left and right side (in a distance d from the square) of

the intersecting line can be varied. Assuming that x  y and x; y  a, nd an expression for

the enclosed area A(x; y) with respect to x and y.



1
Expert's answer
2021-12-07T11:39:44-0500


in "\Delta AKF:"


"tan\theta=\frac{FK}{AK}=\frac{EF-KE}{BC+CD+DE}=\frac{y-x}{\alpha+a+\alpha}"


"tan\theta=\frac{y-x}{a+2\alpha}"



"in\ \Delta AJG:"


"tan\theta=\frac{GJ}{AJ}=\frac{GJ}{\alpha+a}"


"GJ=(a+\alpha)tan\theta=\frac{(y-x)(a+\alpha)}{a+2\alpha}"


"DG=GJ+JD=(a+\alpha)tan\theta+x"



"in\ AID"


"tan\theta=\frac{HI}{AI}=\frac{HI}{\alpha}"


"HI=\alpha tan\theta"


"CH=IC+HI=x+\alpha tan\theta"


enclosed area is a shape of trapezoidal AC//CD


"Area=\frac{1}{2}(CH+DG)\times CD"


"=\frac{1}{2}(x+\alpha tan\theta+(a+\alpha)tan\theta+x)\times a"


"=\frac{1}{2}(2x+(a+\alpha)tan\theta)\times a"


"=\frac{1}{2}(2x+(a+2\alpha)\frac{y-x}{(a+2\alpha)})\times a"


"=\frac{1}{2}(2x+y-x)a"


"=\frac{1}{2}(x+y)a"


"A(x,y)=\frac{a}{2}(x+y)"